# Non-Gaussian Simultaneous Autoregressive Models with Missing Data

### v1

@@ Abstract
Standard simultaneous autoregressive (SAR) models are usually assumed to have normally distributed errors, an assumption that is often violated in real-world datasets, which are frequently found to exhibit non-normal, skewed, and heavy-tailed characteristics. New SAR models are proposed to capture these non-Gaussian features. In this project, the spatial error model (SEM), a widely used SAR-type model, is considered. Three novel SEMs are introduced that extend the standard Gaussian SEM by incorporating Student’s $t$-distributed errors after a one-to-one transformation is applied to the response variable. Variational Bayes (VB) estimation methods are developed for these models, and the framework is further extended to handle missing response data. Standard variational Bayes (VB) methods perform well with complete datasets; however, handling missing data requires a Hybrid VB (HVB) approach, which integrates a Markov chain Monte Carlo (MCMC) sampler to generate missing values. The proposed VB methods are evaluated using both simulated and real-world datasets, demonstrating their robustness and effectiveness in dealing with non-normal data and missing data in spatial models. Although the method is demonstrated using SAR models, the proposed model specifications and estimation approaches are widely applicable to various types of models for handling non-Gaussian data with missing values.
Keywords: Spatial error models; Student’s $t$ errors; Yeo and Johnson (YJ) transformation; Missing not at random; Variational Bayes approximation
@@ 1 Introduction
Conventional spatial statistical models, such as simultaneous autoregressive (SAR) models, conditional autoregressive (CAR) models, and kriging models, among others (see Besag, (1974); Cressie, (1993); Anselin, (1988)), typically assume that the data being modeled follow a Gaussian distribution. However, in practice, this assumption is often violated. In this study, we propose novel SAR-type models and their estimation methods, which are particularly useful when the data are non-Gaussian.
SAR models are a broad class of spatial statistical models commonly used for analysing spatially correlated lattice data (Cressie,, 1993). Examples of SAR models includes spatial error models (SEMs), spatial autoregressive models (SAMs), and spatial Durbin models (SDMs), and many more; see Anselin, (1988), and Cressie, (1993). These models account for spatial dependence, making them useful tools for understanding spatial phenomena in fields such as ecology (Ver Hoef et al.,, 2018), social sciences (Chi and Zhu,, 2008, 2019), finance (Calabrese et al.,, 2017; Agosto et al.,, 2019).
@@
The estimation of SAR models using methods such as the maximum likelihood estimation (MLE) (Ord,, 1975; LeSage,, 1997) and the Bayesian methods (De Oliveira and Song,, 2008; Wu,, 2018), under the assumption of Gaussian errors, has been extensively discussed in the literature. To address the issue of outliers in estimating SEM, Yildirim and Mert Kantar, (2020) proposed a robust version of the ML estimator. Estimation methods for SAR models that do not rely on strong distributional assumptions have been developed. For example, Lee, (2004) introduced the quasi-maximum likelihood estimator (QMLE) for the SAM. Furthermore, Generalised Method of Moments (GMM) estimators, which require fewer distributional assumptions than traditional ML estimators, are robust to unknown heteroskedasticity and deviations from normality. These estimators have been proposed by Kelejian and Prucha, (1998), Lee and Liu, (2010), and Breitung and Wigger, (2018). Non-parametric estimation approaches have also been discussed in the literature for modeling SEM with non-Gaussian data (Al-Momani et al.,, 2017).
@@
Missing values frequently occur in spatial datasets. When estimating SAR models, ignoring missing response values can lead to inconsistency and bias  (Wang and Lee,, 2013; Benedetti et al.,, 2020). Various estimation methods for SAR models with randomly missing data have been well developed (see LeSage and Pace, (2004); Wang and Lee, (2013); Suesse and Zammit-Mangion, (2017); Suesse, (2018); Wijayawardhana et al., (2024)). However, research on estimating SAR models with non-randomly missing data remains relatively limited. Flores-Lagunes and Schnier, (2012) introduced a GMM estimator, while Rabovič and Čížek, (2023) proposed a partial ML method for SEM. Furthermore, Doğan and Taşpinar, (2018) and Seya et al., (2021) used the Metropolis-Hastings (MH) algorithm, which becomes computationally expensive when dealing with a large number of observations.
Variational Bayes (VB) methods have recently emerged as a faster alternative to Markov chain Monte Carlo (MCMC) for estimating complex statistical models (Chappell et al.,, 2008; Han et al.,, 2016; Ong et al.,, 2018; Gunawan et al.,, 2021, 2024). Several commonly used VB methods include mean-field variational Bayes (MFVB) (Ormerod and Wand,, 2010), integrated non-factorised variational Bayes inference (INFVB) (Han et al.,, 2013), and Gaussian variational approximation (Ong et al.,, 2018; Tan and Nott,, 2018). Although VB methods offer a more computationally efficient alternative to MCMC, their use in estimating SAR models has been relatively limited. Wu, (2018) applied two VB approaches, a hybrid MFVB method and an INFVB method, to estimate spatial autoregressive confused (SAC) and matrix exponential spatial specification (MESS) models. Similarly, Bansal et al., (2021) used MFVB and INFVB to estimate spatial count data models, incorporating a MESS model to capture spatial dependence in error terms. More recently, Wijayawardhana et al., (2025) introduced novel VB methods for estimating SAR models with missing data, demonstrating their effectiveness for moderately large datasets with a high proportion of missing values.
@@
Our article makes several key contributions. First, we introduce novel SAR models designed to capture non-Gaussian characteristics of the response variable, such as skewness and heavy tails. These models extend the standard Gaussian SAR by incorporating Student’s t-distributed errors and applying a one-to-one transformation, the Yeo-Johnson (YJ) transformation (Yeo and Johnson,, 2000), to the response variable (see Section 2.2 for details). By combining the properties of the Student’s $t$ distribution and the YJ transformation, these models effectively account for heavy-tailedness and skewness, making them well-suited for complex real-world data. Although we illustrate these new model specifications using SEMs, the framework is applicable to any SAR model.
Second, we propose efficient VB methods for estimating these new SEMs, with and without missing values in the response variable. For the missing data case, we specifically consider the missing not at random (MNAR) mechanism, as defined by Rubin, (1976), where the probability of a missing value depends on both observed and unobserved responses. In standard VB inference, the joint posterior distribution of missing values and model parameters is approximated using a Gaussian variational approximation (Wijayawardhana et al.,, 2025). However, it has been shown that the Gaussian variational approximation with a factor covariance structure (Ong et al.,, 2018), is not sufficiently flexible to jointly approximate the posterior distributions of the SAR model parameters and missing values. Therefore, in this study, we use the standard Gaussian variational approximation of Ong et al., (2018) for SEMs without missing values, while SEMs with missing values are estimated using an extended version of the hybrid VB (HVB) algorithm of Wijayawardhana et al., (2025).
@@
The rest of this paper is organised as follows. Section 2 presents the new SEMs. In Section 3, we present the VB methods to estimate the SEMs with missing data. In Section 4, simulation studies are conducted to evaluate the performance of the new SEMs and VB methods. Section 5 applies the SEMs to a real-world dataset. Section 6 discusses our major results and findings. The paper also has an online supplement with additional technical details and examples.
@@ 2 Simultaneous Autoregressive Models
This section introduces our proposed SAR models, focusing on the spatial error model (SEM), a widely used SAR-type model. We present three novel SEMs that extend the standard Gaussian SEM.
Let $\textbf{y}^{\ast}=(y^{\ast}_{1},y^{\ast}_{2},...,y^{\ast}_{n})^{\top}$ be the vector of response variable observed at $n$ spatial locations $s_{1},\ldots,s_{n}$, X be the $n\times(r+1)$ design matrix containing the covariates, and W be the $n\times n$ spatial weight matrix. The SEM is given by
@@
where e is an $n\times 1$ vector of error terms, $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{r})^{\top}$ is an $(r+1)\times 1$ vector of fixed effects parameters, $\rho$ is the spatial autocorrelation parameter, which measures the strength and direction of spatial dependence (LeSage and Pace,, 2009), and $\textbf{I}_{n}$ is the $n\times n$ identity matrix.
Section 2.1 discusses SEMs with Gaussian and Student’s $t$ errors. Section 2.2 discusses two SEMs with Yeo-Johnson (Yeo and Johnson,, 2000) transformation. Section 2.3 discusses SEMs with missing data.
2.1 Spatial error models with Gaussian and Student’s $t$ errors
The standard SEM assumes Gaussian errors $\textbf{e}\sim N(0,\sigma_{\textbf{e}}^{2}\textbf{I}_{n})$, where $\sigma_{\textbf{e}}^{2}$ is the variance parameter. This assumption implies that the distribution of $\textbf{y}^{\ast}$ is also Gaussian, with the mean vector and the covariance matrix given in Table 1. We call this model the SEM-Gau.
Student’s $t$-distribution has been widely used to model heavy-tailed data in popular statistical models, such as regression (Lange et al.,, 1989; Fernández and Steel,, 1999) and mixed-effects models (Pinheiro et al.,, 2001; Wang and Lin,, 2014). Huang et al., (2021) applied Student’s $t$ errors to spatial autoregressive scalar-on-function regression models. We now discuss SEM-t, which replaces the Gaussian errors in the SEM-Gau model with a Student’s t-distribution. Let $e_{i}\sim t_{\nu}(0,\sigma^{2}_{e})$, where $e_{i}$ is the $i^{th}$ element of the error vector e in Equation (1) and $t_{\nu}(0,\sigma^{2}_{e})$ is the Student’s t distribution with $\nu$ degrees of freedom ($\nu\geq 3$), mean zero, and scale parameter $\sigma^{2}_{e}$. We write Student’s t errors, $e_{i}$, $i=1,...,n$, as a scale mixture of normals (Chan et al.,, 2019),
@@ 2.1 Spatial error models with Gaussian and Student’s $t$ errors
where $IG(a,b)$ denotes the inverse gamma distribution with shape parameter $a$ and scale parameter $b$. The conditional distribution of errors presented in Equation (2), can be written in vector notation as
@@
where the matrix $\boldsymbol{\Sigma}_{\boldsymbol{\tau}}=\textrm{diag}(\tau_{1},\ldots,\tau_{n})$. We now incorporate the error structure from Equation (3) into the SEM described in Equation (1). As a result, the distribution of $\textbf{y}^{\ast}$ given the latent vector $\boldsymbol{\tau}$ is also Gaussian, with the mean vector and the covariance matrix given in Table 1. We refer to this model as SEM-t. The log-likelihood function of $\textbf{y}^{\ast}$ for SEM-Gau and SEM-t is expressed as:
@@
where for the SEM-Gau model, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the SEM-t model, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$. The definitions for r, M, and $\boldsymbol{\phi}$ are given in Table 1.
2.2 Spatial error models with transformations
This section extends the SEM-Gau and SEM-t defined in Section 2.1. We form a novel class of SEMs using an element-wise Yeo and Johnson (YJ) transformation (Yeo and Johnson,, 2000) for the response variable, $\textbf{y}^{\ast}$, in both the SEM-Gau and SEM-t models. We assume that the $i^{th}$ observed response variable $(y_{i})$ is modeled as,
${y}_{i}=t_{{\gamma}}^{-1}(y^{\ast}_{i}),$
where $t_{{\gamma}}^{-1}(.)$ is the inverse function of the YJ transformation. For $20<\gamma<$ 2, it is defined as:
$y_{i}=t_{{\gamma}}^{-1}(y^{\ast}_{i})=\begin{cases}(y^{\ast}_{i}\gamma+1)^{1/\gamma}-1&\text{if }y^{\ast}_{i}\geq 0\\ 1-(-(2-\gamma)y^{\ast}_{i}+1)^{1/(2-\gamma)}&\text{if }y^{\ast}_{i}<0.\\ \end{cases}$ (5)
As the YJ transformation makes the data more symmetric and less skewed, applying its inverse transformation moves the data away from symmetry. This implies that $\textbf{y}=t_{{\gamma}}^{-1}(\textbf{y}^{\ast})=(t^{-1}_{{\gamma}}(y^{\ast}_{1}),t^{-1}_{{\gamma}}(y^{\ast}_{2}),\dots,t^{-1}_{{\gamma}}(y^{\ast}_{n}))^{\top}$ becomes an asymmetric, non-Gaussian response variable. The SEM with the YJ transformation (YJ-SEM) is defined as:
$\textbf{y}=t_{{\gamma}}^{-1}(\textbf{y}^{\ast})=t_{{\gamma}}^{-1}(\textbf{X}\boldsymbol{\beta}+(\textbf{I}_{n}-\rho\textbf{W})^{-1}\textbf{e}).\$ (6)
The density of y for the YJ-SEM is:

### v2

# Bayesian Inference for Non-Gaussian Simultaneous Autoregressive Models with Missing Data

@@ Abstract
Standard simultaneous autoregressive (SAR) models typically assume normally distributed errors, an assumption often violated in real-world datasets that frequently exhibit non-normal, skewed, or heavy-tailed characteristics. New SAR models are proposed to capture these non-Gaussian features. The spatial error model (SEM), a widely used SAR-type model, is considered. Three novel SEMs are introduced, extending the standard Gaussian SEM. These extensions incorporate Student’s $t$-distributed errors to accommodate heavy-tailed behaviour, one-to-one transformations of the response variable to address skewness, or a combination of both. Variational Bayes (VB) estimation methods are developed for these models, and the framework is further extended to handle missing response data under the missing not at random (MNAR) mechanism. Standard VB methods perform well with complete datasets; however, handling missing data requires a hybrid VB (HVB) approach, which integrates a Markov chain Monte Carlo (MCMC) sampler to generate missing values. The proposed VB methods are evaluated using both simulated and real-world datasets, demonstrating their robustness and effectiveness in dealing with non-Gaussian data and missing data in spatial models. Although the method is demonstrated using SAR models, the proposed model specifications and estimation approaches are widely applicable to various types of models for handling non-Gaussian data with missing values.
Keywords: Spatial error models; Student’s $t$ errors; Yeo and Johnson (YJ) transformation; Missing not at random; Variational Bayes
@@ 1 Introduction
Simultaneous autoregressive (SAR) models are a broad class of spatial statistical models commonly used for analysing spatially correlated lattice data, with applications spanning ecology, social sciences, and finance (Cressie,, 1993; Chi and Zhu,, 2008; Calabrese et al.,, 2017; Ver Hoef et al.,, 2018). Examples of SAR models include spatial error model (SEM), spatial autoregressive models (SAM), and spatial Durbin model (SDM), among others (Anselin,, 1988; Cressie,, 1993). Similar to other conventional spatial models, SAR models typically assume Gaussian-distributed data; see  Besag, (1974); Anselin, (1988); Cressie, (1993). However, this normality assumption is frequently violated in practice, as real-world datasets often exhibit non-normal characteristics such as skewness and heavy-tailed distributions.
@@
Several studies have examined the estimation of SAR models when the underlying error distributions are non-Gaussian, leading to methods that do not rely on strong distributional assumptions. For example, quasi-maximum likelihood estimators (QMLE) have been proposed for the SAM (Lee,, 2004), while various Generalised Method of Moments (GMM) approaches have been developed for different types of SAR models, which are robust to unknown heteroskedasticity and deviations from Gaussianity (Kelejian and Prucha,, 1998; Lee and Liu,, 2010; Breitung and Wigger,, 2018). However, these approaches require the response variable to be fully observed and therefore cannot accommodate missing data.
@@
Missing values frequently arise in spatial datasets. When estimating SAR models, ignoring missing response values can lead to inconsistency and bias (Wang and Lee,, 2013; Benedetti et al.,, 2020). Various estimation methods for SAR models with randomly missing data have been well developed (see LeSage and Pace, (2004); Wang and Lee, (2013); Suesse and Zammit-Mangion, (2017); Suesse, (2018); Wijayawardhana et al., (2024)). However, research on estimating SAR models when data are missing not at random (MNAR) remains relatively scarce. Flores-Lagunes and Schnier, (2012) proposed a GMM estimator for this setting, while Doğan and Taşpinar, (2018) and Seya et al., (2021) employed the Metropolis–Hastings (MH) algorithm.
Bayesian estimation of SAR models using Markov chain Monte Carlo (MCMC) methods can be computationally demanding, particularly when the data contain missing values (Wijayawardhana et al.,, 2025). Variational Bayes (VB) provides a computationally efficient alternative, yet it remains underutilised in the context of SAR models. For instance, Wu, (2018) and Bansal et al., (2021) applied VB to fully observed SAR models, while Wijayawardhana et al., (2025) proposed hybrid VB (HVB) approaches for SAR models with missing data. Moreover, Bayesian inference methods, including both VB and MCMC, offer full posterior uncertainty quantification, in contrast to GMM, which yields only point estimates with asymptotic standard errors.
@@
Our article makes several key contributions aimed at addressing limitations in the current literature on SAR models, particularly in handling non-Gaussian responses, missing data, and the computational challenges arising from these complex scenarios. First, we introduce novel SAR models designed to capture non-Gaussian characteristics of the response variable, such as skewness and heavy tails. We extend conventional Gaussian SAR models in two ways: by incorporating Student’s $t$-distributed errors to address heavy-tailedness, and by applying the Yeo-Johnson (YJ) transformation (Yeo and Johnson,, 2000) to the response variable to handle skewness (see Section 2.2). Unlike the traditional two-step approach, where skewness is first reduced using a transformation such as YJ or Box-Cox and then a model is fitted, our integrated framework combines these steps within the SAR structure. These innovations yield three novel SAR models that jointly capture skewness and heavy-tailedness, making them well-suited for complex real-world data. While we demonstrate these models using SEM, yielding three SEM variants, the framework is broadly applicable to other SAR models.
Second, we propose efficient VB methods for estimating these new SEMs, both with and without missing values in the response variable. For SEMs with missing values, we use an extended version of the HVB algorithm of Wijayawardhana et al., (2025) to handle non-Gaussian response variables and missing data under the MNAR mechanism. While the proposed HVB algorithm is an approximation compared to the MCMC methods (Doğan and Taşpinar,, 2018; Seya et al.,, 2021) and Hamiltonian Monte Carlo (HMC) (Neal,, 2011), it offers substantial computational advantages. Our simulation study in Section 6 shows that the HVB algorithm is more than ten times faster than the HMC implementation in RStan (Stan Development Team,, 2023) for a dataset of $n=625$ observations with approximately $50\%$ missing responses. Furthermore, Sections 7 and 8 show that the proposed HVB approach scales efficiently to much larger datasets, for which HMC becomes computationally impractical.
@@
The rest of the paper is organised as follows. Section 2 introduces the new SEMs, and Section 3 discusses SEMs with missing responses. In Section 4, we present the VB methods for estimating SEMs with missing data. Section 5 discusses Bayesian model comparison. Sections 6 and 7 present two simulation studies, and Section 8 applies the SEMs to a real-world dataset. Section 9 summarises the main findings and outlines directions for future work. The paper also has an online supplement with additional technical details. The R code to reproduce all simulations and real-world analyses is available at https://github.com/AnjanaWijayawardhana/Non-Gaussian-SAR-with-Missing-Data-VB.
@@ 2 Simultaneous Autoregressive Models
In this section, we use our proposed framework to extend conventional Gaussian SAR models to non-Gaussian settings. We showcase the extension using spatial error models (SEMs), a widely adopted form of SAR-type models.
Let $\textbf{y}^{\ast}=(y^{\ast}_{1},y^{\ast}_{2},...,y^{\ast}_{n})^{\top}$ be the $n\times 1$ vector of response variable observed at $n$ spatial locations $\textbf{s}_{1},\ldots,\textbf{s}_{n}$, X be the $n\times(r+1)$ design matrix containing the covariates, and W be the $n\times n$ spatial weight matrix. The SEM is given by
@@
where e is the $n\times 1$ vector of error terms, $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{r})^{\top}$ is the $(r+1)\times 1$ vector of fixed effects parameters, $\rho$ is the spatial autocorrelation parameter, which measures the strength and direction of spatial dependence (LeSage and Pace,, 2009), and $\textbf{I}_{n}$ is the $n\times n$ identity matrix. Section 2.1 discusses SEMs with Gaussian and Student’s $t$ errors. Section 2.2 discusses two SEMs with the Yeo-Johnson transformation (Yeo and Johnson,, 2000).
2.1 SEMs with Gaussian and Student’s $t$ errors
The conventional SEM assumes Gaussian error terms $\textbf{e}\sim N(\textbf{0},\sigma_{\textbf{e}}^{2}\textbf{I}_{n})$ in the model presented in Equation (1), where $\sigma_{\textbf{e}}^{2}$ is the error variance parameter. This assumption implies that the distribution of $\textbf{y}^{\ast}$ is also Gaussian, with the mean vector $\textbf{X}\boldsymbol{\beta}$ and covariance matrix $\sigma^{2}_{\textbf{e}}(\textbf{A}^{\top}\textbf{A})^{-1}$, where $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$. We call this model the SEM-Gau.
We now discuss SEM-t, which assumes that the error terms e in Equation (1) follow a Student’s $t$-distribution. Student’s $t$-distribution has been widely used to model heavy-tailed data in popular statistical models, such as regression (Lange et al.,, 1989; Fernández and Steel,, 1999) and mixed-effects models (Pinheiro et al.,, 2001; Wang and Lin,, 2014). Huang et al., (2021) applied Student’s $t$ errors to spatial autoregressive scalar-on-function regression models.
Let $e_{i}\sim t_{\nu}(0,\sigma^{2}_{e})$, where $e_{i}$ is the $i^{\textrm{th}}$ element of the error vector e in Equation (1) and $t_{\nu}(0,\sigma^{2}_{e})$ is the Student’s $t$ distribution with $\nu$ degrees of freedom ($\nu\geq 3$), mean zero, and scale parameter $\sigma^{2}_{e}$. These Student’s $t$ errors, $e_{i}$ for $i=1,\ldots,n$, can be expressed as a scale mixture of normals (Chan et al.,, 2019):
@@ 2.1 SEMs with Gaussian and Student’s $t$ errors
where $IG(a,b)$ denotes the inverse gamma distribution with shape parameter $a$ and rate parameter $b$. The conditional distribution of errors presented in Equation (2), can be written in vector notation as
@@
where the matrix $\boldsymbol{\Sigma}_{\boldsymbol{\tau}}=\textrm{diag}(\tau_{1},\ldots,\tau_{n})$. We now incorporate the error structure from Equation (3) into the SEM described in Equation (1). As a result, the distribution of $\textbf{y}^{\ast}$ conditional on the latent vector $\boldsymbol{\tau}$ is also Gaussian, with the mean vector $\textbf{X}\boldsymbol{\beta}$ and covariance matrix $\sigma^{2}_{\textbf{e}}(\textbf{A}^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A})^{-1}$, where $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$. This model is referred to as the SEM-t. The log-likelihood function of $\textbf{y}^{\ast}$ for both the SEM-Gau and SEM-t is given by:
@@
where for the SEM-Gau, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the SEM-t, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$. The definitions for r, M, and $\boldsymbol{\phi}$ are given in Table 1.
Table 1: The definitions of r, M, and the parameters $\boldsymbol{\phi}$ for SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t. For SEM-Gau and SEM-t, the density of the response vector conditional on $\boldsymbol{\xi}$ follows a multivariate Gaussian distribution with mean $\textbf{X}\boldsymbol{\beta}$ and covariance matrix $\sigma^{2}_{\textbf{e}}\textbf{M}^{-1}$, where $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$.
Model $\boldsymbol{\phi}$ r M
SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\textbf{A}$
SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$
YJ-SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e},},\rho,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\textbf{A}$
YJ-SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$
2.2 SEMs with transformations
This section extends the SEM-Gau and SEM-t in Equation (1) by applying the YJ transformation element-wise to the response vector $\textbf{y}^{\ast}$.
Let $y_{i}$ be the $i^{\textrm{th}}$ element of the response vector in the SEM with YJ transformation. We assume that $y_{i}$ is modelled as
${y}_{i}=t_{{\gamma}}^{-1}(y^{\ast}_{i}),\quad i=1,...,n,$ (5)
where $y^{\ast}_{i}$ is the $i^{\textrm{th}}$ element of the response vector in the SEM given in Equation (1), with either Gaussian or Student’s $t$-distributed errors, and $t_{{\gamma}}^{-1}(\cdot)$ is the inverse function of the YJ transformation. For $0<\gamma<2$, it is defined as:
[…]
As the YJ transformation makes the data more symmetric and less skewed, applying its inverse transformation moves the data away from symmetry. This implies that $\textbf{y}=t_{{\gamma}}^{-1}(\textbf{y}^{\ast})=(t^{-1}_{{\gamma}}(y^{\ast}_{1}),t^{-1}_{{\gamma}}(y^{\ast}_{2}),\dots,t^{-1}_{{\gamma}}(y^{\ast}_{n}))^{\top}$ becomes an asymmetric, non-Gaussian response variable. In vector notation, the SEM with the YJ transformation (YJ-SEM) is defined as:
$\textbf{y}=t_{{\gamma}}^{-1}(\textbf{y}^{\ast})=t_{{\gamma}}^{-1}(\textbf{X}\boldsymbol{\beta}+(\textbf{I}_{n}-\rho\textbf{W})^{-1}\textbf{e}).\\$ (7)
Using the standard technique for transforming random variables, it is straightforward to show that the density of y for the YJ-SEM is: